OpenStudy (anonymous):

The base of a solid is bounded by the graphs of y=3x, y=6, and x=0. The cross sections perpindcular to the x axis are rectangles with perimeter 20, find the volume

OpenStudy (amistre64):



OpenStudy (amistre64):

what does this part tell us: The cross sections perpindcular to the x axis are rectangles with perimeter 20

OpenStudy (anonymous):

yes

OpenStudy (amistre64):

perimeter never changes is what the means to me; P = 2z + 2y; where we know y = 6-3x

OpenStudy (amistre64):

20 = 2z + 2(6-3x) 10 = z+ 6-3x 4+3x = z

OpenStudy (anonymous):

if yes exactly

OpenStudy (amistre64):

so x is from 0 to 2 if i see it right; y = 6 when 3x = 6

OpenStudy (anonymous):

what is it that i am integrating though, i missed that

OpenStudy (anonymous):

like we said z was height right?? and we said z is 4+3x right?

OpenStudy (amistre64):